OpenStudy (anonymous):

If 4(3x-8) = 25. what is 3x-10? How to solve this?

OpenStudy (anonymous):

@DeadPooI

OpenStudy (hazelbrown):

first expand the brackets

OpenStudy (deadpooi):

Hello there Young man.

OpenStudy (deadpooi):

12x - 32 = 25 12x = 57 x = 57 / 12 3(57/12) - 10 = 4.25

OpenStudy (deadpooi):

You need to expand the brackets by multiplying 4 in 3x - 8

OpenStudy (deadpooi):

Click best response if helped you buddy

OpenStudy (anonymous):

I have four options, 8, 23, 11, 25. What now?

OpenStudy (anonymous):

@DeadPooI